# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 14.12 Type ${\bf B}_{2}$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,4$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\bf Sp}_{4}/{B})$. In other words, the above collection is obtained from the collection (79) by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection (85) is right dual to (79).
Proof.
According to Proposition (7.5), one needs to check the following orthogonalities:
  1. [1)]
  2. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-4],\mathcal{L}(-\rho))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-4],{{\sf A}_{i}})=0$ for $i\neq-4$
  3. 2.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  4. 3.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  5. 4.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  6. 5.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  7. 6.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\alpha}})=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  8. 7.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\alpha}})=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  9. 8.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{L}(0))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),{{\sf A}_{i}})=0$ for $i\neq 0$.
[…]
2)2)2 ): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$. We have
$\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\otimes(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\rho)[3]).$ (86)
By Remark 14.34, (3), the latter group is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-\omega_{\beta})\otimes(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\rho)[3])$. Tensoring the sequence (57) with $\mathcal{L}(-\omega_{\beta}-\rho)$ and using Theorem 14.1, wee see from the long cohomology sequence that the required isomorphism follows from $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-2\rho))={\sf k}[-4]$.
[…]
4)4)4 ): $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}$. By Remark 14.34, (1) and (2), ${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})=\mathcal{U}_{2}$ and ${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})=\Psi_{2}^{\omega_{\alpha}}$. The bundle $\Psi_{2}^{\omega_{\alpha}}$ is the cokernel in the short exact sequence
$0\rightarrow\mathcal{L}(-2\omega_{\alpha})\rightarrow\nabla_{\omega_{\alpha}}^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow 0;$ (87)
applying $\mathop{\mathrm{Hom}}\nolimits_{X}(\mathcal{U}_{2}\otimes\mathcal{L}(-\rho),-)$ to this sequence and using Theorem 14.1, we see that the only non-vanishing cohomology group that appears in the long cohomology sequences is $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-4\omega_{\alpha}))={\sf k}[-3]$ (note that this isomorphism holds for $p=2$, despite the restriction on $p$ in Theorem 14.1: there is an isomorphism $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-4\omega_{\alpha}))=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{\beta},\mathcal{L}(-4\omega_{\alpha}))$ and $Y_{\beta}$ is isomorphic to ${\mathbb{P}}(\nabla_{\omega_{\alpha}})$, while the sheaf $\mathcal{L}(-4\omega_{\alpha})$ is the canonical sheaf on $Y_{\beta}$. Therefore, the required isomorphism holds by Serre duality on $Y_{\beta}$). The coboundary homomorphism makes it isomorphic to the group $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho),{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))={\sf k}[-2]$. This cancels out with the shift by $[2]$ giving the required isomorphism. ∎
@@ 14.14 Type ${\bf G}_{2}$
Let $p\geq 5$ be a prime. Let $G$ be a simple group of type ${G}_{2}$. Let $\alpha$ and $\beta$ be the simple roots for ${G}_{2}$ with $\beta$ being the long root. Let $\omega_{\alpha}$ and $\omega_{\beta}$ be the two fundamental weights. Let ${P}_{\alpha}$ (resp., ${P}_{\beta}$) be the minimal parabolic generated by ${B}$ and ${\bf U}_{\alpha}$ (resp., ${B}$ and ${\bf U}_{\beta}$). The Grassmannian ${G}/{P}_{\beta}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})$ via the line bundle $\mathcal{L}(\omega_{\alpha})$ and ${\rm dim}\nabla_{\omega_{\alpha}}=7$. The Grassmannian ${G}/{P}_{\alpha}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\beta}}^{\ast})$ via the line bundle $\mathcal{L}(\omega_{\beta})$ and ${\rm dim}\nabla_{\omega_{\beta}}=14$. As usual, denote $\pi_{\alpha}:{G}/{B}\rightarrow{G}/{P}_{\alpha}$ (resp., $\pi_{\beta}:{G}/{B}\rightarrow{G}/{P}_{\beta}$) the projection. Associated to $\pi_{\alpha}$ and $\pi_{\beta}$ are Demazure operators ${\sf D}_{\alpha}:=\pi_{\alpha}^{\ast}{\pi_{\alpha}}_{\ast}$ and ${\sf D}_{\beta}:=\pi_{\beta}^{\ast}{\pi_{\beta}}_{\ast}$. Associated to the fundamental weights $\omega_{\alpha}$ and $\omega_{\beta}$ are the vector bundles $\Psi_{1}^{\omega_{\alpha}}$ and $\Psi_{1}^{\omega_{\beta}}$ defined as in the sequence (56).
@@
The following collection is a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({G}_{2}/{B})$:
@@
$0\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow{\sf Ker}(\nabla_{\omega_{\alpha}}^{\otimes 2}\rightarrow\nabla_{2\omega_{\alpha}})\otimes\mathcal{O}_{{G}_{2}/{B}}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\alpha})\rightarrow\mathcal{L}(2\omega_{\alpha})\rightarrow 0$ (89)
@@
The case of ${G}_{2}$ differs from other types in rank two in that the bundle $\Psi_{2}^{\omega_{\alpha}}$ appears in the second block of the exceptional collection for ${G}_{2}$ in Lemma 14.37 (as opposed to ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ suggested by the cases of $A_{2}$ and ${B}_{2}$). The reason is that the bundle ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is not exceptional (one of the two deviations in Proposition 14.25): ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is the kernel of the surjective map $\nabla_{\omega_{\beta}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{\beta}\mathcal{L}(\omega_{\beta})$ and the latter vector bundle ${\sf D}_{\beta}\mathcal{L}(\omega_{\beta})$ is not exceptional. In turn, ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is not exceptional either.
@@
The $G$-equivariant vector bundles in the above collection can be explicitly identified:
@@
Let us give concrete descriptions of the high rank $G$-equivariant vector bundles in the above list. The rank two bundle $\mathcal{U}_{2}$ fits into a short exact sequence
@@
the corresponding extension is given by the isomorphism ${\mathrm{Ext}}^{1}_{G/B}(\mathcal{L}(\omega_{\alpha}-\omega_{\beta}),\mathcal{L}(-\omega_{\alpha}))={\rm H}^{1}(G/B,\mathcal{L}(-\beta))=k$.
@@
$0\rightarrow\sf D_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\rightarrow{\sf Ker}(\nabla_{\omega_{\alpha}}^{\otimes 2}\rightarrow\nabla_{2\omega_{\alpha}})\otimes\mathcal{O}_{{G}_{2}/{B}}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{U}_{2}^{\ast}\rightarrow{\sf S}^{2}\mathcal{U}_{2}^{\ast}\rightarrow 0$ (93)
Now $\sf D_{\alpha}(\Psi_{2}^{\omega_{\alpha}})=:\mathcal{E}$ is a vector bundle on $G_{2}/B$ since ${\rm R}^{1}{\pi_{\alpha}}_{\ast}\Psi_{2}^{\omega_{\alpha}}=0$; the latter fact follows from the short sequence
$0\rightarrow\Omega^{2}_{\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})}\otimes\mathcal{L}(2\omega_{\alpha})|_{G/P_{\beta}}\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow\mathcal{L}(0)\rightarrow 0,$ (94)
and from the splitting type of $\Omega^{2}_{\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})}\otimes\mathcal{L}(2\omega_{\alpha})|_{G/P_{\beta}}$ to a line $l$ on $G/P_{\beta}$ (cf. [45, Section 4]) as $(\underbrace{-1,\dots,-1,}_{5}\underbrace{0,\dots,0}_{10})$. It also follows from the above that $\mathcal{E}$ fits into the short exact sequence

### v3

@@ 14.12 Type ${\bf B}_{2}$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,4$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\bf Sp}_{4}/{\mathbf{B}})$. In other words, the above collection is obtained from the collection (87) by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection (91) is right dual to (87).
@@ 14.14 Type ${\bf G}_{2}$
Let $p\geq 5$ be a prime. Let $\mathbf{G}$ be a simple group of type ${\mathbf{G}}_{2}$. Let $\alpha$ and $\beta$ be the simple roots for ${\mathbf{G}}_{2}$ with $\beta$ being the long root. Let $\omega_{\alpha}$ and $\omega_{\beta}$ be the two fundamental weights. Let ${\mathbf{P}}_{\alpha}$ (resp., ${\mathbf{P}}_{\beta}$) be the minimal parabolic generated by ${\mathbf{B}}$ and ${\mathbf{P}}_{\alpha}$ (resp., ${\mathbf{B}}$ and ${\mathbf{P}}_{\beta}$). The Grassmannian ${\mathbf{G}}/{\mathbf{P}}_{\beta}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})$ via the line bundle $\mathcal{L}(\omega_{\alpha})$ and ${\rm dim}\nabla_{\omega_{\alpha}}=7$. The Grassmannian ${\mathbf{G}}/{\mathbf{P}}_{\alpha}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\beta}}^{\ast})$ via the line bundle $\mathcal{L}(\omega_{\beta})$ and ${\rm dim}\nabla_{\omega_{\beta}}=14$. As usual, denote $\pi_{\alpha}:{\mathbf{G}}/{\mathbf{B}}\rightarrow{\mathbf{G}}/{\mathbf{P}}_{\alpha}$ (resp., $\pi_{\beta}:{\mathbf{G}}/{\mathbf{B}}\rightarrow{\mathbf{G}}/{\mathbf{P}}_{\beta}$) the projection. Associated to $\pi_{\alpha}$ and $\pi_{\beta}$ are Demazure operators ${\sf D}_{\alpha}:=\pi_{\alpha}^{\ast}{\pi_{\alpha}}_{\ast}$ and ${\sf D}_{\beta}:=\pi_{\beta}^{\ast}{\pi_{\beta}}_{\ast}$. Associated to the fundamental weights $\omega_{\alpha}$ and $\omega_{\beta}$ are the vector bundles $\Psi_{1}^{\omega_{\alpha}}$ and $\Psi_{1}^{\omega_{\beta}}$ defined as in the sequence (64).
@@
The following collection is a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathbf{G}}_{2}/{\mathbf{B}})$:
@@
$0\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow{\sf Ker}(\nabla_{\omega_{\alpha}}^{\otimes 2}\rightarrow\nabla_{2\omega_{\alpha}})\otimes\mathcal{O}_{{\mathbf{G}}_{2}/{\mathbf{B}}}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\alpha})\rightarrow\mathcal{L}(2\omega_{\alpha})\rightarrow 0$ (93)
@@
The case of ${\mathbf{G}}_{2}$ differs from other types in rank two in that the bundle $\Psi_{2}^{\omega_{\alpha}}$ appears in the second block of the exceptional collection for ${\mathbf{G}}_{2}$ in Lemma 14.38 (as opposed to ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ suggested by the cases of $A_{2}$ and ${\mathbf{B}}_{2}$). The reason is that the bundle ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is not exceptional (one of the two deviations in Proposition 14.26): ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is the kernel of the surjective map $\nabla_{\omega_{\beta}}\otimes\mathcal{L}(0)\rightarrow{\sf D}_{\beta}\mathcal{L}(\omega_{\beta})$ and the latter vector bundle ${\sf D}_{\beta}\mathcal{L}(\omega_{\beta})$ is not exceptional. In turn, ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}$ is not exceptional either.
@@
The $\mathbf{G}$-equivariant vector bundles in the above collection can be explicitly identified:
@@
Let us give concrete descriptions of the high rank $\mathbf{G}$-equivariant vector bundles in the above list. The rank two bundle $\mathcal{U}_{2}$ fits into a short exact sequence
@@
the corresponding extension is given by the isomorphism ${\mathrm{Ext}}^{1}_{\mathbf{G}/\mathbf{B}}(\mathcal{L}(\omega_{\alpha}-\omega_{\beta}),\mathcal{L}(-\omega_{\alpha}))={\rm H}^{1}(\mathbf{G}/\mathbf{B},\mathcal{L}(-\beta))=k$.
@@
$0\rightarrow\sf D_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\rightarrow{\sf Ker}(\nabla_{\omega_{\alpha}}^{\otimes 2}\rightarrow\nabla_{2\omega_{\alpha}})\otimes\mathcal{O}_{{\mathbf{G}}_{2}/{\mathbf{B}}}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{U}_{2}^{\ast}\rightarrow{\sf S}^{2}\mathcal{U}_{2}^{\ast}\rightarrow 0$ (97)
Now $\sf D_{\alpha}(\Psi_{2}^{\omega_{\alpha}})=:\mathcal{E}$ is a vector bundle on $\mathbf{G}_{2}/\mathbf{B}$ since ${\rm R}^{1}{\pi_{\alpha}}_{\ast}\Psi_{2}^{\omega_{\alpha}}=0$; the latter fact follows from the short sequence
$0\rightarrow\Omega^{2}_{\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})}\otimes\mathcal{L}(2\omega_{\alpha})|_{\mathbf{G}/P_{\beta}}\rightarrow\Psi_{2}^{\omega_{\alpha}}\rightarrow\mathcal{L}(0)\rightarrow 0,$ (98)
and from the splitting type of $\Omega^{2}_{\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})}\otimes\mathcal{L}(2\omega_{\alpha})|_{\mathbf{G}/P_{\beta}}$ to a line $l$ on $\mathbf{G}/P_{\beta}$ (cf. [47, Section 4]) as $(\underbrace{-1,\dots,-1,}_{5}\underbrace{0,\dots,0}_{10})$. It also follows from the above that $\mathcal{E}$ fits into the short exact sequence